A 60 – g rifle bullet leaves the muzzle with a velocity of 600 m/s. The 3.0 kg rifle is held firmly by a 60.0 – kg man. With what velocity will the man and rifle recoil?



1
Expert's answer
2021-11-11T17:08:03-0500

Let's apply the law of conservation of momentum:


"m_bv_b+(m_{man}+m_{rifle})v_{recoil}=0,""(m_{man}+m_{rifle})v_{recoil}=-m_bv_b,""v_{recoil}=-\dfrac{m_bv_b}{m_{man}+m_{rifle}},""v_{recoil}=-\dfrac{0.06\ kg\times600\ \dfrac{m}{s}}{60\ kg+3\ kg}=-0.57\ \dfrac{m}{s}."

The sign minus means that the recoil velocity of the man and rifle directed in the opposite direction to the motion of the bullet.